# Accelerating Data Chunking in Deduplication Systems using Vector Instructions

### v1

@@ 6.3. Throughput breakdown - Extreme Byte Search vs Range Scan
Figure 12. Throughput Breakdown with AVX-512 instructions
@@
The throughput impact of each processing pattern depends on algorithmic and dataset characteristics. Figure 12 shows the individual impact of accelerating Extreme Byte Search and Range Scan in VRAM on DEB and LNX with an 8KB chunk size. VRAM-EBS represents RAM running with only Extreme Byte Search acceleration, while VRAM-512 uses both accelerated patterns. Similarly, VMAXP-EBS represents MAXP running with only Extreme Byte Search acceleration, while VMAXP-512 uses both accelerated patterns. We use the DEB and LNX datasets for this experiment.
Figure 12(a) shows that on DEB, VRAM-EBS achieves a throughput of $18.5$ GB/s compared to RAM at $1.7$ GB/s. Accelerating Range Scan provides an additional speedup of $11.4$ GB/s. On the other hand on LNX, VRAM-EBS only achieves $2.7$ GB/s compared to RAM at $2$ GB/s/. Accelerating Range Scan provides an additional speedup of $27.6$ GB/s. Thus, each pattern has a balanced impact on VRAM’s throughput on DEB while Range Scan majorly contributes to throughput on LNX, showing that dataset characteristics impact the throughput breakdown.
@@ 6.4. VectorCDC across different vector instruction sets
SS-CDC (sscdc, ) requires CPUs with vector instruction sets supporting scatter/gather instructions (§3.2). Such CPUs are present only in a small percentage of datacenter nodes today. However, almost all datacenter machines that are currently deployed support other vector instruction sets without this capability, such as SSE-128, AVX-256, NEON-128, and VSX-128 (§2.3).
@@
Figure 14. Accelerating hashless algorithms with VectorCDC across processor architectures, on DEB at an 8KB average chunk size. Note the different y-axis scale on Figures 14(d) and 14(e).
While §4 discusses VectorCDC’s design using AVX-512 instructions, the same methods can be applied to any vector instruction set that supports VCMP, VMAX, and VMASK operations. In this section, we evaluate VectorCDC’s performance with other such vector instruction sets. Figure 14 shows the throughput achieved by AE-Max, AE-Min, MAXP, and RAM accelerated using VectorCDC with SSE-128, AVX-256, NEON-128, and VSX-128 instruction sets. We ran this experiment using the DEB dataset and an average chunk size of 8 KB.
In summary, VectorCDC is compatible with a large range of platforms that support vector instructions while retaining its benefits over hash-based algorithms, unlike SS-CDC (sscdc, ) which requires CPUs with scatter/gather instruction support. While VectorCDC’s speedups depend on the underlying CPU’s vector instruction support, it remains effective on platforms alternative to x86, such as ARM and IBM Power.
@@ 6.4.2. Intel Emerald Rapids and Skylake
On both platforms, all algorithms also benefit from increasing vector widths; that is, higher vector widths lead to higher throughput. The only exception is MAXP, which does not gain as much as the other algorithms with AVX-512 instructions over AVX-256. This is related to the small window sizes used by MAXP for its Extreme Byte Search phases, which do not benefit from high vector widths. However, MAXP still achieves $4.7\times$ and $5.39\times$ speedups with AVX-512 instructions over its unaccelerated version, on the Skylake and Emerald Rapids machines, respectively.
@@ 6.4.3. ARM v8 Atlas
However, MAXP and RAM still achieve $1.93\times$ and $5.32\times$ speedups, respectively, showing that VectorCDC remains beneficial on ARM platforms with NEON-128 support. Note that these numbers are expected to improve in ARM platforms supporting SVE/SVE2 instructions (arm_sve, ), as they offer native VMASK support. However, as we could not obtain such a platform for our evaluation, we leave a detailed SVE/SVE2 performance review as future work.
@@ 7. Related Work
Many efforts have been made to optimize data chunking. MUCH (much, ) and P-Dedupe (p_dedupe, ) use multiple threads to accelerate chunking. RapidCDC (rapidcdc, ) sometimes skips data chunking by predicting the next chunk boundary based on historical data, but requires maintaining additional metadata. Bimodal Chunking (kruus2010bimodal, ) initially splits the data into large chunks, and then divides duplicate adjacent chunks into smaller ones, to enhance space savings. VectorCDC is compatible with all of these approaches, as they build on top of existing CDC algorithms.
Previous work (lowentropy_cloud, ) which analyzes the characteristics of chunks generated by CDC algorithms, is orthogonal to VectorCDC, as vector acceleration does not affect generated chunks.
@@
Several other efforts exist to optimize the other phases of the deduplication pipeline. StoreGPU (storegpu, ) and GPU-Dedup (gpu_dedup, ) accelerate chunk hash computation using GPUs. SiLo (Silo, ), Sparse Indexing (sparse_indexing, ) and Extreme Binning (extremeBinning, ) optimize hash indexing. HYDRAStor (hydrastor, ) is a distributed deduplication system that focuses on data placement. Several studies incorporate delta compression after deduplication to further compress similar but non-duplicate chunks (shilane2012wan, ; zou2022building, ; zhang2023loopdelta, ). These efforts are orthogonal to ours as we accelerate the data chunking phase.
@@
Vector instructions have been widely used to accelerate other storage systems. MinervaFS (minervafs, ) accelerates the computation of transform and basis functions in generalized deduplication with AVX instructions. ICID (inlinededup_nvme, ) records memory-copy operations in a B-Tree for fine-grained deduplication, accelerating tree searches with AVX instructions. AVX-512 conflict detection instructions have been used to accelerate lightweight data compression algorithms (avx_rle, ). Numerous works attempt to accelerate collision-resistant hashing algorithms used across storage systems with vector instructions (clhash_vectorized, ; murmur_vectorized, ). These efforts are orthogonal to ours as we focus on using vector instructions to accelerate CDC algorithms for block-level deduplication.
@@
Several efforts build end-to-end deduplication systems for encrypted data (secure_dedup_survey, ). They mainly target encryption schemes (mle, ; secure_asokan, ) for the underlying data or focus on reducing attacks on the system (side_channel_dedup, ; sgxdedup, ). Some target specific applications, such as distributing encrypted docker images  (sun2024simenc, ) and encrypted videos (securededup_video, ). As all of these efforts layer encryption atop existing data chunking algorithms, VectorCDC is compatible with all these approaches.
@@ 8. Conclusion
We present VectorCDC, a methodology for accelerating content-defined chunking using vector instructions. VectorCDC avoids the pitfalls of previous work that accelerates CDC algorithms by choosing hashless CDC instead. VectorCDC accelerates these algorithms using novel tree-based search and packed scanning methods. Our evaluation shows that VectorCDC achieves $8.35\times$-$26.2\times$ higher throughput than existing AVX-based CDC techniques. We have made our code publicly available by integrating it with DedupBench (dedupbench, ), and hosted one of our datasets on Kaggle (deb_dataset, ).
@@ Acknowledgements.
We thank the anonymous USENIX FAST 2025 and ACM Transactions on Storage reviewers for their feedback. We thank Lori Paniak for his technical assistance throughout the project, and Mu’men Al-Jarah for his feedback on an earlier version of this work. The research team was supported by grants from the National Cybersecurity Consortium (NCC), Natural Sciences and Engineering Research Council of Canada (NSERC), and the Ontario Research Fund’s Research Excellence Program (ALLRP-561423-20, RGPIN-2025-03332, and ORF-RE012-051). The team was also supported by research grants from Acronis, Oracle Research Labs, and Rogers Communications. Sreeharsha is supported by the Cheriton Graduate Scholarship and the Ontario Graduate Scholarship.

### v2

@@ 6.3. Throughput breakdown - Extreme Byte Search vs Range Scan
Figure 12. Throughput Breakdown with AVX-512 instructions. Note that VRAM-EBS and VMAXP-EBS represent VRAM and VMAXP with only Extreme Byte Search accelerated.
@@
The throughput impact of each processing pattern depends on algorithmic and dataset characteristics. Figure 12 shows the individual impact of accelerating Extreme Byte Search and Range Scan using VRAM on the DEB and LNX datasets with an 8KB chunk size. VRAM-EBS and VMAXP-EBS represent RAM and MAXP running with only Extreme Byte Search acceleration, while VRAM-512 and VMAXP-512 use both accelerated patterns.
Figure 12(a) shows that on DEB, VRAM-EBS achieves a throughput of $18.5$ GB/s compared to RAM at $1.7$ GB/s. Accelerating Range Scan provides an additional speedup of $11.4$ GB/s. On the other hand on LNX, VRAM-EBS only achieves $2.7$ GB/s compared to RAM at $2$ GB/s. Accelerating Range Scan provides an additional speedup of $27.6$ GB/s. Thus, each pattern has a balanced impact on VRAM’s throughput on DEB, while Range Scan primarily contributes to throughput on LNX, indicating that dataset characteristics affect the throughput breakdown.
@@ 6.4. VectorCDC across different vector instruction sets
VectorCDC is compatible with a large range of platforms that support vector instructions such as SSE-128, AVX-256, NEON-128, and VSX-128 (§2.3). This is unlike SS-CDC (Ni et al., 2019) which requires CPUs with scatter/gather instruction support. Such CPUs are present only in a small percentage of datacenter nodes today.
@@
Figure 14. Accelerating hashless algorithms with VectorCDC across processor architectures, on DEB at an 8KB average chunk size. Data labels show speedups over the respective native algorithm for the specific cases discussed in text. Note the different y-axis scale on Figures 14(d) and 14(e).
While §4 discusses VectorCDC’s design using AVX-512 instructions, the same methods can be applied to any vector instruction set that supports VCMP, VMAX, and VMASK operations. In this section, we evaluate VectorCDC’s performance with other such vector instruction sets. We ran this experiment using the DEB dataset and an average chunk size of 8 KB.
@@ 6.4.2. Intel Emerald Rapids and Skylake
On both platforms, all algorithms also benefit from increasing vector widths; that is, higher vector widths lead to higher throughput. The only exception is MAXP, which does not gain as much as the other algorithms with AVX-512 instructions over AVX-256. This is related to the small window sizes used by MAXP for its Extreme Byte Search phases, which do not benefit from high vector widths. However, MAXP still achieves $4.7\times$ and $5.42\times$ speedups with AVX-512 instructions over its unaccelerated version, on the Skylake and Emerald Rapids machines, respectively.
@@ 6.4.3. ARM v8 Atlas
However, MAXP and RAM still achieve $1.93\times$ and $5.32\times$ speedups, respectively, showing that VectorCDC remains beneficial on ARM platforms with NEON-128 support. Note that these numbers are expected to improve in ARM platforms supporting SVE/SVE2 instructions (Shi et al., 2025), as they offer native VMASK support. However, as we could not obtain such a platform for our evaluation, we leave a detailed SVE/SVE2 performance review as future work.
@@ 6.4.4. IBM Power 8
6.5. Evaluation Summary
To summarize, the main takeaways from our evaluation are the following:
  • VectorCDC-based hashless algorithms achieve $15.3\times$$207.2\times$ and $8.35\times$$26.2\times$ higher throughput than unaccelerated and vector-accelerated hash-based algorithms respectively, showing that hashless algorithms are better candidates for vector acceleration (§6.2).
  • VectorCDC effectively alleviates the data chunking performance bottleneck in the deduplication pipeline (§6.2.3).
  • Accelerating both Extreme Byte Search and Range Scan is important because their individual impact depends on dataset and algorithmic characteristics (§6.3).
  • VectorCDC provides benefits across different processor architectures, and is compatible with a wide range of vector instruction sets (§6.4).
  • Accelerating hashless algorithms with VectorCDC does not impact their space savings and generates chunks identical to their unaccelerated counterparts.
  • Hashless algorithms achieve space savings values comparable to or better than those of their hash-based counterparts on real-world datasets (§6.1). The best performing hashless algorithm varies by dataset, showing that accelerating all of them is equally important.
@@ 7. Related Work
Many efforts have been made to optimize data chunking. MUCH (Won et al., 2015) and P-Dedupe (Xia et al., 2012) use multiple threads to accelerate chunking. RapidCDC (Ni and Jiang, 2019) sometimes skips data chunking by predicting the next chunk boundary based on historical data, but requires maintaining additional metadata. Bimodal Chunking (Kruus et al., 2010) initially splits the data into large chunks, and then divides duplicate adjacent chunks into smaller ones, to enhance space savings. VectorCDC is compatible with all of these approaches, as they build on top of existing CDC algorithms.
Previous work (Jarah et al., 2024) that analyzes the characteristics of chunks generated by CDC algorithms, is orthogonal to VectorCDC, as vector acceleration does not affect generated chunks.
Our previous paper at USENIX FAST 2025 (Udayashankar et al., 2025a) presented VectorCDC’s design, but does not discuss accelerating MAXP (Bjørner et al., 2010) or VectorCDC’s performance on varying CPU architectures. Additionally, it does not present a comprehensive evaluation of VectorCDC’s capabilities.
@@
Several other efforts exist to optimize the other phases of the deduplication pipeline. StoreGPU (Al-Kiswany et al., 2008) and GPU-Dedup (Suttisirikul and Uthayopas, 2012) accelerate chunk hash computation using GPUs. SiLo (Xia et al., 2015), Sparse Indexing (Lillibridge et al., 2009) and Extreme Binning (Bhagwat et al., 2009) optimize hash indexing. HYDRAStor (Dubnicki et al., 2009) is a distributed deduplication system that focuses on data placement. Several studies incorporate delta compression after deduplication to further compress similar but non-duplicate chunks (Shilane et al., 2012a; Zou et al., 2022; Zhang et al., 2023). These efforts are orthogonal to ours as we accelerate the data chunking phase.
@@
Vector instructions have been widely used to accelerate other storage systems. MinervaFS (Nielsen et al., 2021) accelerates the computation of transform and basis functions in generalized deduplication with AVX instructions. ICID (Liu et al., 2024) records memory-copy operations in a B-Tree for fine-grained deduplication, accelerating tree searches with AVX instructions. AVX-512 conflict detection instructions have been used to accelerate lightweight data compression algorithms (Ungethum et al., 2018). Numerous works attempt to accelerate collision-resistant hashing algorithms used across storage systems with vector instructions (Lemire and Kaser, 2016; Pan et al., 2018). These efforts are orthogonal to ours as we focus on using vector instructions to accelerate CDC algorithms for block-level deduplication.
@@
Several efforts build end-to-end deduplication systems for encrypted data (Shin et al., 2017). They mainly target encryption schemes (Bellare et al., 2013; Liu et al., 2015) for the underlying data or focus on reducing attacks on the system (Harnik et al., 2010; Ren et al., 2021). Some target specific applications, such as distributing encrypted docker images  (Sun et al., 2024) and encrypted videos (Zheng et al., 2017). As all of these efforts layer encryption atop existing data chunking algorithms, VectorCDC is compatible with all these approaches.
@@ 8. Conclusion
We present VectorCDC, a methodology for accelerating content-defined chunking using vector instructions. VectorCDC avoids the pitfalls of previous work that accelerates CDC algorithms by choosing hashless CDC algorithms instead. VectorCDC accelerates these algorithms using novel tree-based search and packed scanning methods. Our evaluation shows that VectorCDC achieves $8.35\times$-$26.2\times$ higher throughput than existing vector-accelerated CDC algorithms and $15.3\times$-$207.2\times$ higher throughput than unaccelerated algorithms. We have made our code publicly available by integrating it with DedupBench (Liu et al., 2023), and published one of our datasets on Kaggle (Udayashankar et al., 2025b).
@@ Acknowledgements.
We thank the anonymous reviewers of USENIX FAST 2025 and ACM Transactions on Storage for their feedback. We thank Lori Paniak for his technical assistance throughout the project, and Mu’men Al-Jarah for his feedback on an earlier version of this work. The research team was supported by grants from the National Cybersecurity Consortium (NCC), Natural Sciences and Engineering Research Council of Canada (NSERC), and the Ontario Research Fund’s Research Excellence Program (ALLRP-561423-20, RGPIN-2025-03332, and ORF-RE012-051). The team was also supported by research grants from Acronis, Oracle Research Labs, and Rogers Communications. Sreeharsha is supported by the Cheriton Graduate Scholarship and the Ontario Graduate Scholarship.